Two stars of equal mass M are in orbit around one another. They are a distance R apart. What is the gravitational field strength at the point A, midway between them?


(a) GM/R^2

(b) zero

(c) 2GM/R^2

(d) 4GM/R^2




1
Expert's answer
2021-02-18T18:53:23-0500
"g(R)=\frac{GM}{R^2}-\frac{GM}{R^2}=0"

Answer: (b) zero